OpenStudy (anonymous):

find the common ratio of the following geometric sequence. 5/12, 1/3, 4/15, 16/75, 64/375 if necessary use the slash bar to enter a fraction. reduce fractions to their lowest terms

OpenStudy (anonymous):

The common ratio is basically the ratio of a term to a previous term. So the common term of this one is: (1/3)/(5/12) = 4/5

OpenStudy (anonymous):

You can prove this by checking the other terms: (4/15)/(1/3)=(16/75)/(4/15)=(64/375)/(16/75)=4/5 so the common ratio is 4/5

OpenStudy (anonymous):

Refer to the calculation attached using Mathematica.